OpenStudy (konradzuse):

8.3 #2b Find *(T[3] o T[2], o) T[1] (x, y) b. > T[1](x, y) = (x+y, y, -x); T[2](x, y, z) = (0, x+y+z, 3*y); T[3](x, y, z) = (3*x+2*y, 4*z-x-3*y);

OpenStudy (konradzuse):

@tkhunny

OpenStudy (konradzuse):

I got first we do T3 o T2 3*x+6*y, 4*z-x-9*y; 3 x + 6 y, 4 z - x - 9 y > -3*x+6*y, 4*z+x-9*y;

OpenStudy (konradzuse):

and sorry for formatting :P

OpenStudy (tkhunny):

Notation Thoughts: f(g(x)) = f o g This would seem to suggest that you need to start from the RIGHT! This may be a surprise. T[1](x, y) = (x+y, y, -x) T[2](x+y, y, -x) = (0, 2y, 3*y) You do the last one.

OpenStudy (konradzuse):

hmmm....

OpenStudy (konradzuse):

The way it looked in the book and previous answers that you basically take the values of T1 and pujt them in t2 for example the answer in the book was t1(x,y) =( 2x,3y), T2(x,y,) = (x-y,x+y) The answer was 2x-3y, 2x+3y

OpenStudy (konradzuse):

@Kainui